Operations Report – January 26th


Name of person filing report: Carlos Salazar

Non-nominal systems: none

Notes on non-nominal systems: none

Spirit rover used: No
Hours: 153.7
Beginning charge: 100%
Ending charge: 100%
Currently charging: yes

Opportunity rover used: No
Hours: 82.9
Beginning charge: 100%
Ending charge: 100%
Currently charging: Yes

Curiosity rover used: No
Hours: 165.4
Beginning charge: 100%
Ending charge: 100%
Currently charging: yes

Perseverance rover used: no
Hours: 226.5
Beginning charge: 100%
Ending charge: 100%
Currently charging: Yes

General notes and comments: none

Summary of Hab operations:
WATER USE: 28 gallons
Water (static tank): 277 gallons
Water (loft tank): 32 gallons
Water Meter: currently 0155133.5
Static to Loft Pump used – yes
Static tank pipe heater (on or off): on
Static tank heater (on or off) on

Toilet tank emptied: yes

Summary of internet: nominal

Summary of suits and radios: nominal

Summary of GreenHab operations:

WATER USE: 8 gallons
Temperature: 90°F
Heater: On
Supplemental light: Yes
Harvest: 9g Chamomile, 3g Mejorana

Summary of Science Dome operations:
SOC: 100%

Summary of RAM operations: none

Summary of any observatory issues: none

Summary of health and safety issues: none

Questions, concerns and requests to Mission Support:

White flies are accumulating in the plants of the green-hab, especially in the red cherry tomatoes. The insect killer is almost over, is there any possibility to resupply?

Additionally, Felipe, the crew scientist measured the T° inside of the green-hab with a device that he brought and it appeared to be more than 20° F higher than what the sensor inside of the green-hab showed. He also performed a T° measurement with the same device in the main hab and the T° appeared to be the same of the sensor inside of the main hab, which means that is possible that the sensor of the green-hab is not working correctly which worries the green-hab officer a lot because the T° in the green-hab in the day rises up to 90-100°F, and the plants could be receiving much more heat than they actually need. For instance, today the T° that he measured in the green-hab was 134°F while the T° in the sensor inside was 103°F. María Paula says she is opening the door to lower the T° to 90° approximately.

EVA Report – January 26th

Crew 226 EVA Report 26-01-2022

EVA #: 7

Author: Maria Bustos, Felipe Torres

Purpose of EVA: Recognition of the lithology and geological description of the zone. Collection of clay samples.

Start time: 10:00 am
End time: 12:20 pm

Narrative: Several kinds of rocks were recognized, mostly sandstones and conglomerates. Also, ripples, wavy stratification, and other sedimentary structures were recognized. All the areas had a lot of signs of weathering, several kinds of regolith were also found, the most weathered formation was the one that contained the clays.

The clays did not show a clear stratification, and fresh clays were not so easy to find. At the end of the EVA, we found the clays in between the regolith. These clays were of white color and had some metallic minerals, we predict they may be pyrite, covelline, or chalcopyrite.

Further studies are necessary to continue analyzing the geology of the zone.
Destination: Phobos Peak

Coordinates (use UTM NAD27 CONUS): 519300, 4250300

EVA Participants: Yael Mendez, Carlos Salazar, Maria Bustos

Road(s) and routes per MDRS Map: MDRS Entrance Road

Mode of travel: by foot

Felipe Torres, Crew Scientist

Journalist Report – January 26th

Journalist Report 26Jan2022
Yael Méndez
Commander- Crew 226

The days on Mars have passed quietly. The entire crew is completing their projects and preparing for their trip back to Earth.

The morning began with a delicious tortilla breakfast with the combination of ingredients that all of us chose (corn, cheese, onions, pepper sauce), and Paula was the one who cooked it.

Later, we prepared for a very productive EVA, at 10 am, heading for Phobos Peak. A walk filled with colors so we got to know the soil and rock conditions of this place. Sand and clay drew the landscapes of Mars.

After collecting samples, we headed to the Hab, where Cristian and Felipe were waiting for us with a delicious spinach soup with pieces of chicken.

Today I thought that each one of us is very important to this crew, we all have something to teach. Today, for example, Carlos gave us an engineering class, he really has the soul of a teacher.

The life experiences of each one have strengthened us as a great team. Science and engineering stories, ideals to form a company in the aerospace field, love relationships experiences of each crew member, and the calm and comfortable way of taking everything as learning, have given us the motivation to continue with this dream to explore the red planet.

Certainly, each one will go their own way and possibly it will be years before we meet again but in my mind, they will always be the memories, the experiences, and the laughter of the Crew 226, my crew.

Sol Summary – January 26th

Sol: 9

Summary Title: Exploring Mars 🔴

Author’s name: Yael Méndez- Commander

Mission Status: A day to carry out tasks on the station and to continue learning about the Geology of Mars.

Sol Activity Summary: We started the day with a tasty egg omelet breakfast.

We visited Phobos Peak to learn about the soil conditions of the area and take some samples of clays and sands.

Look Ahead Plan: Continue working on mission projects and preparing our trip back to Earth.

Anomalies in work: None

Weather: Sunny in the morning and afternoon.

Crew Physical Status: Nominal

EVA: One EVA realized

Reports to be filed: Sol Summary, Operations Report, EVA Report, HSO Report, Daily photos, Journalist report

Support Requested: Dehydrated lime

Copyright © The Mars Society. All rights reserved. | Main Site